OpenStudy (anonymous):

Are the lines y = –x – 4 and 5x + 5y = 20 perpendicular? Explain. Yes; their slopes are equal. Yes; their slopes have product –1. No; their slopes are not equal. No; their slopes are not opposite reciprocals.

OpenStudy (amistre64):

what are their slopes?

OpenStudy (anonymous):

they dont give me that

OpenStudy (amistre64):

the equations give you that information ... which tells me that you are not familiar with the basics of this material ....

OpenStudy (anonymous):

so explain

OpenStudy (anonymous):

If an equation has the form y = mx + c, m is the slope.

OpenStudy (amistre64):

slope is the measure of the change in y values as a ration to the change in the x values ... the most basic way to find the slope is to define x as 0 and get the y value; then define x as 1 and get the y value and see what the difference between the y values are

OpenStudy (amistre64):

y = –x – 4 ; when x=0, what does y =?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

-0-4=

OpenStudy (anonymous):

-4

OpenStudy (amistre64):

good; and: y = –x – 4 ; when x=1?

OpenStudy (anonymous):

-5 :)

OpenStudy (amistre64):

good :) now we want to know the change that happens in y from -4 to -5; -5-(-4) = ?

OpenStudy (anonymous):

-1

OpenStudy (amistre64):

tha tis correct; so our slope of (y = –x – 4) is: -1

OpenStudy (amistre64):

same process with the other equation: 5x + 5y = 20 ; when x=0, y has to equal what?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

because u subtract on both sides

OpenStudy (anonymous):

and then you get 5y=15

OpenStudy (anonymous):

and divide

OpenStudy (amistre64):

good thinking there, just needs some practice to get better; but the process your thinking of should go like this: 5(0) + 5y = 20 0 + 5y = 20 5y = 20 y = ?

OpenStudy (anonymous):

4 :) oh i see what i did wrong

OpenStudy (amistre64):

correct :) what you were thinking was ahead of the curve when x=1 we get: 5(1) +5y = 20 5 +5y = 20 -5 -5 ------------- 0 +5y = 15 5y = 15 y = ?

OpenStudy (anonymous):

3 :)

OpenStudy (amistre64):

so y moves in this equation from 4 to 3 4 + m = 3 m = 3-4 = ?

OpenStudy (anonymous):

? huh? lol it has 2 eqals?

OpenStudy (amistre64):

its just read as m equals 3-4; so 3-4 = ? when everything equals each other like that you can simply write it out as such

OpenStudy (amistre64):

m = 3-4 3-4 = ? m = ?

OpenStudy (anonymous):

m= 3-4

OpenStudy (amistre64):

good; and 3-4 is equal to what? 3 -4 = ?

OpenStudy (anonymous):

m :)

OpenStudy (amistre64):

lol ... yes, but what what does "m" equal when we actually subtract 4 from 3?

OpenStudy (anonymous):

ooooh lol 1 lol

OpenStudy (amistre64):

3-4 = -1 ; all assume that was a typo :) so if you recall what we started with, finding slopes we determined that y=-x -4 ; has a change in y values of -1; it has a slope of -1 5x + 5y = 20; has a change in y values of -1; it too has a slope of -1 when 2 lines are going in the same directions; are they parallel or perpendicular?

OpenStudy (anonymous):

they are parallel

OpenStudy (amistre64):

very good :) Since they are parallel, then they are definitely NOT perpendicular Yes; their slopes are equal. <-- NOT the answer Yes; their slopes have product –1. <-- NOT the answer Which leaves us with 2 possibilitites No; their slopes are not equal. No; their slopes are not opposite reciprocals.

OpenStudy (amistre64):

No; their slopes are not equal. <-- NOT the answer because they HAVE equal slopes No; their slopes are not opposite reciprocals. <-- MUST be the answer since its the only one left :)

OpenStudy (anonymous):

rights :) wow this helped alot!!

OpenStudy (amistre64):

im glad it did :)